OpenStudy (anonymous):

Simplify 5 root(4)(256x^3) completely. Write answer in radical form. Assume x>0.

OpenStudy (lgbasallote):

\[\Large 5\sqrt[4]{256x^3}?\]

OpenStudy (lgbasallote):

hint: \[\sqrt[4]{256x^3} \implies \sqrt[4]{4 \times 4 \times 4 \times 4 \times x^3}\] does that help?

OpenStudy (anonymous):

@lgbasallote So thats what its wanting me to do?

OpenStudy (lgbasallote):

depends on how you interpreted what i said..

OpenStudy (anonymous):

so would it be 16xroot(4)(x)

OpenStudy (anonymous):

@lgbasallote

OpenStudy (lgbasallote):

nope..

OpenStudy (lgbasallote):

\[\sqrt[4]{4 \times 4 \times 4 \times 4 \times x^3} \implies \sqrt[4]{4^4 \times x^3} \implies \sqrt[4]{4^4} \times \sqrt[4]{x^3}\]

OpenStudy (anonymous):

Oh I sort it... so its 256/4

OpenStudy (lgbasallote):

is that giving you an idea now?

OpenStudy (anonymous):

Yeah so is it 4xroot(4)(x)

OpenStudy (anonymous):

@lgbasallote

OpenStudy (anonymous):

Am I right now?

OpenStudy (lgbasallote):

still no

OpenStudy (lgbasallote):

\[\Large \sqrt[4]{x^3} \ne x\sqrt[4]{x}\]

OpenStudy (lgbasallote):

\[\sqrt[4]{x^5} = x\sqrt[4]{x}\]

OpenStudy (anonymous):

OMG! This is so frustrating I feel dumb :(

OpenStudy (lgbasallote):

you cant take x^3 out of the 4th root

OpenStudy (anonymous):

I still don't get but you can take 2 and leave one x inside right? what happenes to the 256 I still don't get it? Thats not divided by four

OpenStudy (anonymous):

@lgbasallote

OpenStudy (lgbasallote):

it's not divided by 4

OpenStudy (lgbasallote):

if it's \(\sqrt x^3\) then you're right it's \(x\sqrt x\) however it's 4th root

OpenStudy (lgbasallote):

so you cant do it

OpenStudy (lgbasallote):

\[\sqrt x^3 = x^{3/2} \] \[\sqrt[4]{x^3} = x^{3/4}\] do you see the difference?

OpenStudy (lgbasallote):

i dont get where you're getting confused...will you tell me where?

OpenStudy (lgbasallote):

i think you're too accustomed to square root that you're looking at this like square root. this is 4th root

OpenStudy (anonymous):

So then it cant be done?

OpenStudy (anonymous):

@lgbasallote

OpenStudy (lgbasallote):

perhaps a demonstration \[\large \sqrt[4]{16 x^2} \implies \sqrt[4]{2 \times 2 \times 2 \times 2 \times x^2} \implies \sqrt[4]{2^4 x^2} \implies 2\sqrt[4]{x^2} \implies 2\sqrt x\]

OpenStudy (lgbasallote):

here's another one \[\large \sqrt[4]{81x^3} \implies \sqrt[4]{3 \times 3 \times 3 \times 3 \times 3 \times x^3} \implies \sqrt[4]{3^4 x^3} \implies 3\sqrt[4]{x^3}\]

OpenStudy (anonymous):

so then its root(4)(4^4x^3) 4root(4)(x^3) ? @lgbasallote

OpenStudy (lgbasallote):

another one \[\large\sqrt[4]{256x^5} \implies \sqrt[4]{4 \times 4 \times 4 \times 4 \times x^5} \implies \sqrt[4]{4^4 x^5} \implies 4x\sqrt[4]{x}\]

OpenStudy (lgbasallote):

yes!!!

OpenStudy (lgbasallote):

but wiat..that's just \[\sqrt[4]{256x^3}\] remember the question is \[\large 5\sqrt[4]{256x^3}\]

OpenStudy (lgbasallote):

so then you do \[5 \times 4 \sqrt[4]{x^3}\]

OpenStudy (anonymous):

\[20\root(4)\sqrt{x^3}\]

OpenStudy (anonymous):

20root(4)(X^3)

OpenStudy (lgbasallote):

right

OpenStudy (anonymous):

WOOOOO THANKS SO MUCHHHHH!!!! :D YOU ARE THE BEST!!! @lgbasallote

OpenStudy (lgbasallote):

haha you're welcome ^_^

OpenStudy (anonymous):

@lgbasallote So how would I do -5root(5)(32x^13)

OpenStudy (lgbasallote):

32 = 2^5

OpenStudy (lgbasallote):

\[\large\sqrt[5]{32x^{13}} \implies \sqrt[5]{2^5 x^{13}}\]

OpenStudy (lgbasallote):

\[\implies \sqrt[5]{2^5 \times x^{10} \times x^3}\]

OpenStudy (lgbasallote):

do you see it now?

OpenStudy (anonymous):

2x^2root(5)(x^3) ? @lgbasallote

OpenStudy (lgbasallote):

yes!!!

OpenStudy (anonymous):

WOOOOOO!!! Now I get it more it's just confusing the sqrt gets me every time @lg

OpenStudy (lgbasallote):

i need to go now so i hope you get these :DDD

OpenStudy (lgbasallote):

haha i know..it just takes practice and experience

OpenStudy (anonymous):

Thanks so m,ouch again! God Bless!!!! @lgbasallote